OpenStudy (anonymous):

find the x-intercept of 4x+5y=25

OpenStudy (amistre64):

one little note is that there is a not used much form called the intercept form; and all you do is divide it all by that 25 to get: \[\frac{4x}{25}+\frac{5y}{25}=1\] \[\frac{x}{25/4}+\frac{y}{25/5}=1\] \[\frac{x}{25/4}+\frac{y}{5}=1\]

OpenStudy (amistre64):

now it easy to tell that when x = 0; y = 5 and when y=0, x=25/4

OpenStudy (amistre64):

yint (0,5) xint(25/4 ,0)

OpenStudy (anonymous):

What are the y and x intercepts for x+5y=25?

OpenStudy (amistre64):

x +5y = 25 ; /25 x/25 +5y/25 = 1 x/25 + y/5 = 1 ; when x=0, y=5; yint = (0,5) x/25 + y/5 = 1 ; when y=0, x=25; xint = (25,0)